OpenStudy (anonymous):

~Medal+Fanning~ Solve the following system. x − y = 10 x + y = 6 (−2, 8) (−8, 2) (8, −2) (8, 2)

OpenStudy (wolf1728):

That seems pretty easy. A) x − y = 10 B) x + y = 6 Add A and B 2x = 16 x = 8 8 + y = 16 y = -2

OpenStudy (wolf1728):

TYPO = should be 8 -2 = 6